OpenStudy (anonymous):

Find the area of a trapezoid with a height of 6 m and bases of 13 m and 26 m

OpenStudy (anonymous):

\[A = \frac{1}{2}(b_{1}+b_{2})h = \frac{1}{2}(13+26)(6)\]

OpenStudy (anonymous):

a = hight * (b1 + b2)/2 a = 6* (13 +26)/2 a = 6 * 19.5 a = 117